Saidhari bought a toaster for $60. The price was 1/4 off the list price. What was the list price?
Answer by checkley77(12844)

60=x-x/4
60=(4x-x)/4
60*4=3x
240=3x
x=240/3
x=$80 was the original price.
Proof:
80(1-1/4)=60
80(3/4)=60
240/4=60
60=60